OpenStudy (anonymous):

Solve each of the systems by graphing y=-3/8x-6 y=1/2x+1

OpenStudy (anonymous):

Okay sam I need your help on this explain all this to me real simple. Without writing just numbers. but do not solve the problem. Get close to solving the problem but don't solve it.

OpenStudy (anonymous):

please help

sam (.sam.):

You can do like this, y=mx+b b will be y-intercept m will be gradient then you need a coordinate to plot it, any coordinate will do, lets say y=3 for y=-3/8x-6, then x=-24 (-24,3)